Please help me, I know this is an easy one but I cannot get the equation. Please translate the problem to an equation. Jeremy is one more than six times as old as Barbara. The difference of their ages is 66. How Old are they?
I have X=Barbara Age
X+1(6)=Jeremy's age X+X+1(6)=66 What is my problem
Thank you

Answer by Paul(988)

x = barbara age
1+6x = Jeremy.

1+6x-x=66
1+5x=66
5x=65
x=13

13(6)+1=79

Hence, Jeremy is 79 years old and Barbara is 13 years old.
Paul.